OpenStudy (anonymous):

An automobile traveling along a straight road increases its speed from 24 ft/s to 37 ft/s in 180 ft. If the acceleration is constant, how much time elapses while the auto moves the 180 ft? Answer in units of s

OpenStudy (kropot72):

The average speed while traveling the 180 ft is given by: \[\large Average\ speed\ while\ accelerating=\frac{24+37}{2}\ ft/s\] The elapsed time while the auto moves the 180 ft is given by the distance divided by the average speed: \[\large Elapsed\ time=\frac{180\times2}{(24+37)}=(you\ can\ calculate)\ s\]